# Client Clustering Meets Knowledge Sharing: Enhancing Privacy and Robustness in Personalized Peer-to-Peer Learning

### v1

@@ Abstract
Index Terms:
peer-to-peer machine learning, decentralized learning, personalization, differential privacy, poisoning attacks.
@@ I Introduction
Fully decentralized or peer-to-peer (P2P) learning has emerged as a promising direction for enabling personalized learning in heterogeneous environments such as those found in IoT ecosystems. In contrast to centralized federated learning (FL), P2P learning eliminates the need for a central server and distributes computation among participating clients [8, 5]. This approach leverages direct communication between clients with their own data distributions, offering potential benefits such as faster convergence [9, 5] and enhanced data privacy compared to centralized FL shi2023improving[10, 11]. In the context of IoT, where devices often operate under bandwidth and power constraints, P2P learning allows edge devices to collaboratively train models by communicating only with a limited number of neighbors, depending on their communication and computational capabilities [12, 13].
Personalized decentralized learning faces three key challenges. The first is clustering, where clients seek similar peers to share knowledge. Clustering algorithms [1, 14] group clients with similar data distributions to train localized models, improving accuracy. However, most methods rely on predefined cluster counts and require a central server [15, 16, 17, 18], which is impractical due to communication bandwidth constraints. More critically, these approaches risk privacy leaks as clients share data for clustering. The second challenge is data privacy—clients’ data often contains sensitive information, requiring strong privacy guarantees. A common solution is adding differential privacy (DP) noise and clipping gradients [19, 20, 21, 22], but this degrades model performance [23, 24, 20]. In P2P settings, this effect is amplified due to fewer model updates (gradients) that are used in model aggregation [25]. Finally, the third challenge is the presence of malicious clients affected by data poisoning [26] or model poisoning [27, 28, 29, 30] attacks, which can degrade the final model performance. Due to its fully decentralized nature, P2P learning is highly vulnerable to poisoning attacks, in which malicious clients can exploit the system by sending carefully crafted local models to their neighboring peers. A common mitigation method is to use recent decentralized FL (DFL) defenses [31, 32, 33] or adapt existing centralized FL ones [34, 35, 36, 37] to DFL. However, a defense that is not tailored to personalized P2P settings may fail in scenarios with highly heterogeneous clients and can introduce additional computational overhead if each client performs robust aggregation individually [38].
@@
In this paper, we propose P4 (Personalized, Private, Peer-to-Peer) [^1] https://anonymous.4open.science/r/private_decentralized_learning-F413, a method designed to deliver personalized models for resource-constrained IoT devices while ensuring differential privacy and robustness against poisoning attacks. In P4, clients first group themselves in a fully decentralized manner using a similarity metric based on their model weights. Once grouped, clients perform collaborative training with upper-bounded data privacy risk and robust aggregation against adversarial model updates. To enhance privacy-utility trade-offs, P4 incorporates the proxy model [47] to facilitate knowledge distillation among clients. This separation between the local and proxy models results in fast convergence and personalized parameter adaptation for each client (as shown in Section V-B), while also enhancing resilience against data poisoning attacks (as discussed in Section IV-E). To mitigate model poisoning attacks, P4 employs a combined defense strategy based on m-Krum [36] and anomaly-detection [46], which utilizes client similarity within each group to filter out malicious updates. P4 is well-suited for IoT deployments that involve numerous clients and non-stationary data distributions, where communication and computational efficiency are critical [48].
@@ III-A System Model
$F_{i}(w_{i})=\mathbb{E}[\mathcal{L}_{(x,y)\sim D_{i}}(w_{i};(x,y))]$ (1)
To find parameter to minimize the expected loss, each client seeks to determine an optimal parameter set, denoted as $w_{i}^{*}$, within the parameter space $w_{i}\in\mathbb{R}^{d}$. This optimization aims to minimize the expected loss over the client’s own local data distribution $\ell(f_{i},y)$, represented in Equation 2:
$w_{i}^{*}=arg\min_{w_{i}\in\mathbb{R}^{d}}\mathbb{E}_{(x,y)\sim D_{i}}\ell(f_{i}(w_{i};x),y)]$ (2)
However, solely using client’s local dataset is not sufficient to find a generalizable parameter for each client, so local training leads to poor generalization performance [5, 7]. In order to achieve good generalization, each client is willing to collaborate and share knowledge (i.e., their model gradients) with a subset of other clients. We note that the goal of the learning is not to reach a ”global consensus” model, but for each client to obtain personalized models that perform well on their own data distributions.
@@ III-B Threat Model
From a security standpoint, we consider adversaries in a semi-honest setting, where clients can deviate from the protocol specification by contributing malicious updates designed to degrade the performance or integrity of the final model. In particular, our threat model assumes the adversary is capable of orchestrating Sybil attacks [74], thereby controlling multiple peers within the network. However, consistent with prior literature [40, 41], we constrain the adversary’s stake to no more than $30\%$ of the total participants in the network. Specifically, our threat model includes prevalent attack scenarios investigated in recent privacy-preserving and peer-to-peer FL frameworks [42, 40, 41, 43]. These include label flipping  [26], byzantine-zero  [27], byzantine-random  [27], and byzantine-flip  [28]. These attacks collectively represent critical and realistic threats faced by distributed machine learning protocols, particularly in peer-to-peer and privacy-preserving contexts. In our label flipping setting, client data is poisoned by flipping all labels as in [26]. Byzantine attacks corrupt local model updates by setting weights to zero, random values, or flipped values using $w_{g}+(w_{g}-w_{l})$, where $w_{g}$ is the global model, and $w_{l}$ is the real local model. Additionally, we assume an attacker may have knowledge of the system design, particularly the group formation phase (see Section IV-B).
@@ IV-A Overview
Figure 1 illustrates the overall design of the P4 approach, which addresses these challenges. It operates in two phases: group formation and co-training. During group formation, clients privately identify peers with similar data distributions and form groups accordingly. In the co-training phase, clients train their local and proxy models and share locally computed data gradients with added noise only within their group. A group aggregator then filters out adversarial model updates, aggregates the remaining gradients, and sends the resulting global model update back to the clients. The following sections provide a more detailed explanation of the system. A table with all the notation is provided in Appendix A
Figure 1: The overall design of the P4 approach. Clients employ local ($\phi$) and proxy ($\theta$) models, aggregating only the proxy model.
@@ IV-B Phase 1: Group Formation
In federated learning, similar data distributions among clients plays an especially important role for model convergence [75, 76, 77]. Selecting groups of clients for co-training without considering the underlying data heterogeneity—such as through random clustering—can lead to performance worse than standalone local training [78, 1]. Therefore, identifying clients with similar data distributions is critical for effective collaborative training.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ IV-C Phase 2: Co-training
To train the local model and proxy model together, inspired by [25], we apply knowledge distillation between these two models [47]. More specifically, training of the local model will be done by classification loss (Equation 8) and KL divergence loss (Equation 9):
$\mathbf{L}_{CE}(f_{\theta_{i}})=\mathbb{E}_{(x,y)\sim D_{i}}CE[f_{\theta_{i}}(x)\parallel y]$ (8)
@@
$\mathbf{L}_{KL}(f_{\theta_{i}};f_{\phi_{i}})=\mathbb{E}_{(x,y)\sim D_{i}}KL[f_{\theta_{i}}(x)\parallel f_{\phi_{i}}(x)]$ (9)
In addition, to distill knowledge from the local model to the proxy model, we use the Kullback–Leibler divergence described in Equation 9 to make the proxy model output closer to the local model output.
Therefore, the proxy model learns from both the local model and the local data by combining the classification loss and the KL divergence loss shown in Equation 10. This way, the proxy model learns to correctly predict the true label of training instances as well as to match the probability estimate of its local model.
$\mathbf{L}_{\theta_{i}}=(1-\alpha)\cdot\mathbf{L}_{CE}(f_{\theta_{i}})+\alpha\cdot\mathbf{L}_{KL}(f_{\theta_{i}};f_{\phi_{i}})$ (10)
@@
The objective of the local model is shown in Equation 11.
$\mathbf{L}_{\phi_{i}}=(1-\beta)\cdot\mathbf{L}_{CE}(f_{\phi_{i}})+\beta\cdot\mathbf{L}_{KL}(f_{\phi_{i}};f_{\theta_{i}})$ (11)
@@
With this architecture, the local model $f_{\phi_{i}}$ can extract personalized model weights for the client without dealing with DP noise, leading to improved performance. In other words, by decoupling local models from the differential privacy noise, we aim to enhance model performance, resulting in models that better align with the specific data characteristics of individual clients. Beside that, the parameters of proxy model $f_{\theta_{i}}$ could be aggregated with neighbor clients to benefit from their knowledge under DP guarantees.
We note that it is not necessary for each member of a group to receive model updates $\Delta\theta_{j}$ from all other members and update its proxy model $\theta_{i}$. Instead, every few rounds, an aggregator can be selected periodically from within the group. Therefore, other clients within the group only send their model updates $\Delta\theta_{j}$ to the aggregator client and subsequently receives the aggregated model $\theta$ from that client, which is used for the next round of co-training.
@@ IV-D Differential Privacy
To ensure privacy during co-training (see Algorithm 2), each gradient of a proxy model of $i$-th client is divided by its norm as described in Equation 12 and then the noise is added to the gradient as shown in Equation 13 (line 6 in Algorithm 2).
$\tilde{g}_{ij}=\frac{g_{ij}}{\max(1,\|g_{ij}\|_{2}/\mathcal{C})}$ (12)
$\tilde{H_{i}}=\frac{1}{sR}\Sigma_{j\in S_{i}}\tilde{g}_{ij}+\frac{2\mathcal{C}}{sR}\mathcal{N}(0,\sigma_{g}^{2})$ (13)
@@
$\sigma_{g}=\Omega(\frac{s\sqrt{lTK\log(\frac{2Tl}{\delta})\log(\frac{2}{\delta})}}{\epsilon\sqrt{M^{\prime}}})$ (14)
@@ IV-E2 Model Poisoning Tolerance
m-Krum defines a Krum aggregation rule $\text{Kr}(w_{i},\ldots,w_{n})$ as follows. For any $i\neq j$, they denote by $i\rightarrow j$ the fact that $w_{j}$ belongs to $n-f-2$ closest vectors to $w_{i}$, where $n$ is the total number of clients and $f$ is the number of malicious clients. Then, they define for each client $i$, the score $s(i)=\sum_{i\rightarrow j}\|w_{i}-w_{j}\|^{2}$, where the sum runs over $n-f-2$ closest vectors to $w_{i}$ based on an Euclidean distance. Finally, $\text{Kr}(w_{i},\ldots,w_{n})=w_{i_{*}}$, where $i_{*}$ refers to the client minimizing the score, $s(i_{*})\leq s(i)$ for all $i$. An $m$ parameter in m-Krum defines the number of clients to choose that have minimum scores.
@@ V-A1 Datasets
Dataset # classes # clients # samples (per client) # features
@@ V-A2 Non-IID Setting
To generate non-IID tasks and analyze the impact of data heterogeneity, we adopt two common methods: (i) alpha-based heterogeneity and (ii) sharding-based heterogeneity. Both partition data among clients to create non-IID distributions. In the alpha-based approach, most of client’s data comes from a single class, whereas in the sharding-based method, each client exclusively receives data from a specific subset of multiple classes. Sharding-based heterogeneity follows the procedure of [5], where a dataset with $L$ classes is divided into $P$ shards per class, generating $M=\frac{LP}{N}$ tasks. Each task consists of $N$ randomly assigned classes, with one shard per class. We evaluate the effect of heterogeneity by setting $N\in\{2,4,8\}$. Alpha-based heterogeneity, used in prior work [24, 90], controls heterogeneity through a parameter $\gamma$. For each client, $\gamma\%$ of the data is sampled IID from all classes, while the remaining $1-\gamma\%$ comes from a single dominant class. We assess heterogeneity effects using $\gamma\in\{25\%,50\%,75\%\}$.
@@ V-B Privacy-Utility Amplification
In our experiments with a CNN model in Figure 3, which possesses more parameters compared to the linear model, previous approaches suffered from noisy training due to differential privacy noise and collaborative training with clients with different data distribution. In contrast, P4 consistently converges to robust parameter solutions across all settings, even with a minimal number of communication rounds.

### v2

@@ I Introduction
Fully decentralized or peer-to-peer (P2P) learning has emerged as a promising direction for enabling personalized learning in heterogeneous environments such as those found in IoT ecosystems. In contrast to centralized federated learning (FL), peer-to-peer learning eliminates the need for a central server and distributes computation among participating clients [8, 5]. This approach leverages direct communication between clients with their own data distributions, offering potential benefits such as faster convergence [9, 5] and enhanced data privacy compared to centralized federated learning shi2023improving[10, 11]. In the context of IoT, where devices often operate under bandwidth and power constraints, P2P learning allows edge devices to collaboratively train models by communicating only with a limited number of neighbors, depending on their communication and computational capabilities [12, 13].
Personalized decentralized learning faces three key challenges. The first is clustering, where clients seek similar peers to share knowledge. Clustering algorithms [1, 14] group clients with similar data distributions to train localized models, improving accuracy. However, most methods rely on predefined cluster counts and require a central server [15, 16, 17, 18], which is impractical due to communication bandwidth constraints. More critically, these approaches risk privacy leaks as clients share data for clustering. The second challenge is data privacy—clients’ data often contains sensitive information, requiring strong privacy guarantees. A common solution is adding differential privacy (DP) noise and clipping gradients [19, 20, 21, 22], but this degrades model performance [23, 24, 20]. In the peer-to-peer settings, this effect is amplified due to fewer model updates (gradients) that are used in model aggregation [25]. Finally, the third challenge is the presence of malicious clients affected by data poisoning [26] or model poisoning [27, 28, 29, 30] attacks, which can degrade the final model performance. Due to its fully decentralized nature, peer-to-peer learning is highly vulnerable to poisoning attacks, in which malicious clients can exploit the system by sending carefully crafted local models to their neighboring peers. A common mitigation strategy is to use recent decentralized FL (DFL) defenses [31, 32, 33] or adapt existing centralized FL ones [34, 35, 36, 37] to DFL. However, a defense that is not tailored to personalized peer-to-peer settings may fail in scenarios with highly heterogeneous clients and can introduce additional computational overhead, particularly when each client is required to perform robust aggregation individually [38].
@@
In this paper, we propose P4 (Personalized, Private, Peer-to-Peer) [^1] https://github.com/mammadmaheri7/private_decentralized_learning, a method designed to deliver personalized models for resource-constrained IoT devices while ensuring differential privacy and robustness against poisoning attacks. In P4, clients first group themselves in a fully decentralized manner using a similarity metric based on their model weights. Once grouped, clients perform collaborative training with upper-bounded data privacy risk and robust aggregation against adversarial model updates. To enhance privacy-utility trade-offs, P4 incorporates the proxy model [47] to facilitate knowledge distillation among clients. This separation between the local and proxy models results in fast convergence and personalized parameter adaptation for each client (as shown in Section V-B), while also enhancing resilience against data poisoning attacks (as discussed in Section IV-E). To mitigate model poisoning attacks, P4 employs a combined defense strategy based on m-Krum [36] and anomaly-detection [46], which utilizes client similarity within each group to filter out malicious updates. P4 is well-suited for IoT deployments that involve numerous clients and non-stationary data distributions, where communication and computational efficiency are critical [48].
@@ III-A System Model
$F_{i}(w_{i})=\mathbb{E}[\mathcal{L}_{(x,y)\sim D_{i}}(w_{i};(x,y))].$ (1)
To find parameters to minimize the expected loss, each client seeks to determine an optimal parameter set, denoted as $w_{i}^{*}$, within the parameter space $w_{i}\in\mathbb{R}^{d}$. This optimization aims to minimize the expected loss over the client’s own local data distribution $\ell(f_{i},y)$, represented in Equation 2:
$w_{i}^{*}=arg\min_{w_{i}\in\mathbb{R}^{d}}\mathbb{E}_{(x,y)\sim D_{i}}\ell(f_{i}(w_{i};x),y)].$ (2)
However, solely using client’s local dataset is not sufficient to find generalizable parameters for each client, so local training leads to poor generalization performance [5, 7]. In order to achieve good generalization, each client is willing to collaborate and share knowledge (i.e., their model gradients) with a subset of other clients. We note that the goal of the learning is not to reach a “global consensus” model, but for each client to obtain personalized models that perform well on their own data distributions.
@@ III-B Threat Model
From a security standpoint, we consider adversaries in a semi-honest setting, where clients can deviate from the protocol specification by contributing malicious updates designed to degrade the performance or integrity of the aggregated model. In particular, our threat model assumes the adversary is capable of orchestrating Sybil attacks [74], thereby controlling multiple peers within the network. However, consistent with prior literature [40, 41], we constrain the adversary’s stake to no more than $30\%$ of the total participants in the network. Specifically, our threat model includes prevalent attack scenarios investigated in recent privacy-preserving and peer-to-peer FL frameworks [42, 40, 41, 43]. These include label flipping  [26], byzantine-zero  [27], byzantine-random  [27], and byzantine-flip  [28]. These attacks collectively represent critical and realistic threats faced by distributed machine learning protocols, particularly in peer-to-peer and privacy-preserving contexts. In our label flipping setting, client data is poisoned by flipping all labels as in [26]. Byzantine attacks corrupt local model updates by setting weights to zero, random values, or flipped values using $w_{g}+(w_{g}-w_{l})$, where $w_{g}$ is the global model, and $w_{l}$ is the real local model. Additionally, we assume an attacker may have knowledge of the system design, particularly the group formation phase (see Section IV-B).
@@ IV-A Overview
Figure 1 illustrates the overall design of the P4 approach, which addresses these challenges. It operates in two phases: group formation and co-training. During group formation, clients privately identify peers with similar data distributions and form groups accordingly. In the co-training phase, clients train their local and proxy models and share locally computed data gradients with added noise only within their group. A group aggregator then filters out adversarial model updates, aggregates the remaining gradients, and sends the resulting global model update back to the clients. The following sections provide a more detailed explanation of the system. A table with all the notation is provided in Appendix A.
@@ IV-B Phase 1: Group Formation
In federated learning, similar data distributions among clients play an especially important role for model convergence [75, 76, 77]. Selecting groups of clients for co-training without considering the underlying data heterogeneity—such as through random clustering—can lead to performance worse than standalone local training [78, 1]. Therefore, identifying clients with similar data distributions is critical for effective collaborative training.
@@
Figure 1: The overall design of the P4 approach. Clients employ local ($\phi$) and proxy ($\theta$) models, aggregating only the proxy model. Groups are formed based on the $\ell_{1}$-norm dissimilarity of their model weights. Within each group, updates are exchanged via one client acting as an aggregator, which may change during training to balance communication overhead. After receiving the aggregated model, clients perform local training for personalization.
@@
$\underset{G}{min}\sum_{i=1}^{M}\sum_{j=1}^{M}\mathbb{I}[G(i,j)=1]\lVert\mathbf{w}_{i}-\mathbf{w}_{j}\rVert_{1},$ (6)
@@
21 if $\text{{similar\_clients}}[i]\neq\emptyset$ then Continue
@@
23 if $i=mcs[k][0][0]$ and $\text{{similar\_clients}}[k]=\emptyset$ then
24    $\text{{similar\_clients}}[i]\leftarrow\text{{similar\_clients}}[i]\cup\{k\}$
@@
26    $\text{{similar\_clients}}[k]\leftarrow\text{{similar\_clients}}[k]\cup\{i\}$
@@ IV-C Phase 2: Co-training
To train the local model and the proxy model together, inspired by [25], we apply knowledge distillation between these two models [47]. More specifically, training of the local model will be done by classification loss (Equation 8) and KL divergence loss (Equation 9):
$\mathbf{L}_{CE}(f_{\theta_{i}})=\mathbb{E}_{(x,y)\sim D_{i}}CE[f_{\theta_{i}}(x)\parallel y],$ (8)
@@
In addition, to distill knowledge from the local model to the proxy model, we use the Kullback–Leibler divergence described in Equation 9 to make the proxy model output closer to the local model output:
$\mathbf{L}_{KL}(f_{\theta_{i}};f_{\phi_{i}})=\mathbb{E}_{(x,y)\sim D_{i}}KL[f_{\theta_{i}}(x)\parallel f_{\phi_{i}}(x)].$ (9)
Therefore, the proxy model learns from both the local model and the local data by combining the classification loss and the KL divergence loss shown in Equation 10. This way, the proxy model learns to correctly predict the true label of training instances as well as to match the probability estimate of its local model:
$\mathbf{L}_{\theta_{i}}=(1-\alpha)\cdot\mathbf{L}_{CE}(f_{\theta_{i}})+\alpha\cdot\mathbf{L}_{KL}(f_{\theta_{i}};f_{\phi_{i}}).$ (10)
@@
The objective of the local model is shown in Equation 11:
$\mathbf{L}_{\phi_{i}}=(1-\beta)\cdot\mathbf{L}_{CE}(f_{\phi_{i}})+\beta\cdot\mathbf{L}_{KL}(f_{\phi_{i}};f_{\theta_{i}}).$ (11)
@@
With this architecture, the local model $f_{\phi_{i}}$ can extract personalized model weights for the client without dealing with DP noise, leading to improved performance. In other words, by decoupling local models from the differential privacy noise, we aim to enhance model performance, resulting in models that better align with the specific data characteristics of individual clients. Besides that, the parameters of the proxy model $f_{\theta_{i}}$ could be aggregated with neighbor clients to benefit from their knowledge under DP guarantees.
We note that it is not necessary for each member of a group to receive model updates $\Delta\theta_{j}$ from all other members and update its proxy model $\theta_{i}$. Instead, every few rounds, an aggregator can be selected periodically from within the group. Therefore, other clients within the group only send their model updates $\Delta\theta_{j}$ to the aggregator client and subsequently receive the aggregated model $\theta$ from that client, which is used for the next round of co-training.
@@ IV-D Differential Privacy
To ensure privacy during co-training (see Algorithm 2), each gradient of a proxy model of $i$-th client is divided by its norm as described in Equation 12 and then the noise is added to the gradient as shown in Equation 13 (line 6 in Algorithm 2):
$\tilde{g}_{ij}=\frac{g_{ij}}{\max(1,\|g_{ij}\|_{2}/\mathcal{C})},$ (12)
$\tilde{H_{i}}=\frac{1}{sR}\Sigma_{j\in S_{i}}\tilde{g}_{ij}+\frac{2\mathcal{C}}{sR}\mathcal{N}(0,\sigma_{g}^{2}).$ (13)
@@
$\sigma_{g}=\Omega(\frac{s\sqrt{lTK\log(\frac{2Tl}{\delta})\log(\frac{2}{\delta})}}{\epsilon\sqrt{M^{\prime}}}).$ (14)
@@ IV-E2 Model Poisoning Tolerance
m-Krum defines a Krum aggregation rule $\text{Kr}(w_{i},\ldots,w_{n})$ as follows. For any $i\neq j$, they denote by $i\rightarrow j$ the fact that $w_{j}$ belongs to nf2n-f-2 closest vectors to $w_{i}$, where $n$ is the total number of clients and $f$ is the number of malicious clients. Then, they define for each client $i$, the score $s(i)=\sum_{i\rightarrow j}\|w_{i}-w_{j}\|^{2}$, where the sum runs over nf2n-f-2 closest vectors to $w_{i}$ based on an Euclidean distance. Finally, $\text{Kr}(w_{i},\ldots,w_{n})=w_{i_{*}}$, where $i_{*}$ refers to the client minimizing the score, $s(i_{*})\leq s(i)$ for all $i$. An $m$ parameter in m-Krum defines the number of clients to choose that have minimum scores.
@@ V-A1 Datasets
Dataset # classes # clients
# samples
(per client)
# features
@@ V-A2 Non-IID Setting
To generate non-IID tasks and analyze the impact of data heterogeneity, we adopt two common methods: (i) alpha-based heterogeneity and (ii) sharding-based heterogeneity. Both partition data among clients to create non-IID distributions. In the alpha-based approach, most of the client’s data comes from a single class, whereas in the sharding-based method, each client exclusively receives data from a specific subset of multiple classes. Sharding-based heterogeneity follows the procedure of [5], where a dataset with $L$ classes is divided into $P$ shards per class, generating $M=\frac{LP}{N}$ tasks. Each task consists of $N$ randomly assigned classes, with one shard per class. We evaluate the effect of heterogeneity by setting $N\in\{2,4,8\}$. Alpha-based heterogeneity, used in prior work [24, 90], controls heterogeneity through a parameter $\gamma$. For each client, $\gamma\%$ of the data is sampled IID from all classes, while the remaining $1-\gamma\%$ comes from a single dominant class. We assess heterogeneity effects using $\gamma\in\{25\%,50\%,75\%\}$.
@@ V-B Privacy-Utility Amplification
In our experiments with a CNN model in Figure 3, which possesses more parameters compared to the linear model, previous approaches suffered from noisy training due to differential privacy noise and collaborative training with clients with different data distributions. In contrast, P4 consistently converges to robust parameter solutions across all settings, even with a minimal number of communication rounds.